OpenStudy (anonymous):

Simplify 3 square root of 5 end root minus 2 square root of 7 end root plus square root of 45 end root minus square root of 28.

OpenStudy (anonymous):

anyone? please i dont understand

jimthompson5910 (jim_thompson5910):

The expression is this right? \[\large 3\sqrt{5}-2\sqrt{7}+\sqrt{45}-\sqrt{28}\]

OpenStudy (anonymous):

do you know how to solve it?

jimthompson5910 (jim_thompson5910):

first we need to simplify \(\Large \sqrt{45}\)

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{45}=\sqrt{9*5}\] \[\large \sqrt{45}=\sqrt{9}*\sqrt{5}\] \[\large \sqrt{45}=3\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

Notice how I factored it into 9*5 one of the factors is the largest perfect square factor possible

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

how would you simplify \(\Large \sqrt{28}\) ?

OpenStudy (anonymous):

2 square root of 7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

So \[\large 3\sqrt{5}-2\sqrt{7}+\sqrt{45}-\sqrt{28}\] turns into \[\large 3\sqrt{5}-2\sqrt{7}+3\sqrt{5}-2\sqrt{7}\]

jimthompson5910 (jim_thompson5910):

from here combine like terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont really know how to do this part its confusing

jimthompson5910 (jim_thompson5910):

if I gave you something like 3x+7y+2x+10y, would you be able to combine like terms?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it would be 5x + 17y right?

jimthompson5910 (jim_thompson5910):

yes that's correct

jimthompson5910 (jim_thompson5910):

you'll use the same idea

jimthompson5910 (jim_thompson5910):

let x = sqrt(5) and y = sqrt(7) \[\large 3\color{red}{\sqrt{5}}-2\color{blue}{\sqrt{7}}+3\color{red}{\sqrt{5}}-2\color{blue}{\sqrt{7}}\] \[\large 3\color{red}{x}-2\color{blue}{y}+3\color{red}{x}-2\color{blue}{y}\]

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

simplify `3x-2y+3x-2y` to get ???

OpenStudy (anonymous):

6x-y?

jimthompson5910 (jim_thompson5910):

6x is correct -y is not

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

what would the answer be?

jimthompson5910 (jim_thompson5910):

-2y-2y is -4y, agreed?

jimthompson5910 (jim_thompson5910):

\[\large 3\color{red}{x}-2\color{blue}{y}+3\color{red}{x}-2\color{blue}{y}\] \[\large 6\color{red}{x}-4\color{blue}{y}\] \[\large 6\color{red}{\sqrt{5}}-4\color{blue}{\sqrt{7}}\]

jimthompson5910 (jim_thompson5910):

So, \[\large 3\sqrt{5}-2\sqrt{7}+\sqrt{45}-\sqrt{28} = 6\sqrt{5}-4\sqrt{7}\]

OpenStudy (anonymous):

ok thank a lot

OpenStudy (anonymous):

can you help me with this one please? if you want to. Which statement is true about the difference 2 square root of 7 − square root of 28?

jimthompson5910 (jim_thompson5910):

\[2\sqrt{7} - \color{red}{\sqrt{28}} = 2\sqrt{7} - \color{red}{2\sqrt{7}} = 0\sqrt{7} = 0\]